# SPAWNing Structural Priming Predictions from a Cognitively Motivated Parser

### v1

@@ Abstract
Structural priming is a widely used psycholinguistic paradigm to study human sentence representations. In this work we propose a framework for using empirical priming patterns to build a theory characterizing the structural representations humans construct when processing sentences. This framework uses a new cognitively motivated parser, SPAWN, to generate quantitative priming predictions from theoretical syntax and evaluate these predictions with empirical human behavior. As a case study, we apply this framework to study reduced relative clause representations in English. We use SPAWN to generate priming predictions from two theoretical accounts which make different assumptions about the structure of relative clauses. We find that the predictions from only one of these theories (Participial-Phase) align with empirical priming patterns, thus highlighting which assumptions about relative clause better capture human sentence representations.
@@ 1 Introduction
Prior work found that targets like 1 were produced more often, and were processed more rapidly, when they were preceded by primes like 1, that have the same structure, than when they were preceded by primes like 1, which, while describing the same transfer event as 1, have a different structure.
@@
From this result, Branigan et al. (1995) inferred that participants’ mental representation of 1 is more similar to that of 1 than of 1 .
Branigan and Pickering (2017) propose that by carefully studying which sentences prime each other we can build a theory of human structural representations. Building such a theory requires us to generate hypotheses about what are the interesting prime-target pairs to compare. Insights from theoretical syntax, a field that has spent decades studying the structure of sentences, can help constrain this hypothesis space Gaston et al. (2017). In this work we introduce a new parser, the Serial Parser in ACT-R With Null elements (SPAWN), that can generate quantitative priming predictions from theories in syntax. By comparing SPAWN predictions from competing theoretical accounts, we can study which theoretical differences result in differing priming predictions, and therefore might be meaningful for psycholinguistics. Then, by comparing these predictions to empirical priming behavior in humans, we can study which theoretical assumptions are consistent with the representations that humans construct when processing sentences.
SPAWN is a cognitively motivated parser in which all of the parsing decisions are based on the computational principles proposed by a general purpose cognitive architecture, Adaptive Control of Thought-Rational (ACT-R; Anderson et al., 2004). Thus, SPAWN not only describes the computations underlying human parsing (Marr’s computational level), but also specifies the cognitive processes involved (Marr’s algorithmic level). This level of specification is necessary to explain why a sentence A is primed more by sentence B compared to C, and therefore necessary to generate quantitative behavioral priming predictions from syntactic theories. Existing algorithmic models of parsing Lewis and Vasishth (2005) can account for only a limited set of linguistic phenomena and theoretical assumptions. SPAWN can model a wider range of phenomena and assumptions because it uses a more flexible grammar formalism and specifies mechanisms to handle covert lexical items.
@@
We generate priming predictions from two competing syntactic theories: Whiz-Deletion (Chomsky, 1965), which assumes that the structure of 1 is identical to the structure of 1, and Participial-Phase (Harwood, 2018) which assumes that 1 and 1 have different structures. We describe these theories in more detail in § 4. We generate six sets of predictions from the two theories by modulating the strength of prior knowledge and the specific re-analysis mechanism used. Then, we compare the predictions from these two theories to empirical human data we collected using a novel web-based comprehension-to-production priming paradigm.
We found that the predictions from the Whiz-Deletion never aligned with the qualitative pattern of human priming behavior, whereas the predictions from Participial-Phase theory aligned with the empirical pattern when the predictions are generated from a SPAWN model with very weak to no prior knowledge; this observation held for the two variations of the reanalysis the mechanisms we tried. Taken together, these results suggest that the Participial-Phase account better characterizes human sentence representations, and more broadly highlight how SPAWN can be used to adjudicate between competing theoretical assumptions.
@@ 2 Background
2.1 ACT-R framework
ACT-R is a cognitive architecture with modules which are designed to explain general cognition through a small set of general computational principles and mechanisms that are relevant to a wide range of tasks and domains. One such mechanism which is particularly relevant in SPAWN is the retrieval of information from memory. The specific computational principles and algorithms that guide retrieval in ACT-R are outlined in § 3.2.1. Crucially, since ACT-R is intended to be a general purpose cognitive mechanism, most of the hyperparameters involved in this algorithm are already fixed based on data from a wide range of experimental paradigms and cognitive phenomena. This is useful because it restricts the degrees of freedom, thereby constraining the space of predictions that can be generated from any given theory.
@@ 2.2 Prior models of parsing
In most existing symbolic and neural network based parsers, the parsing decisions are not driven by specific cognitive principles such as the ones proposed by ACT-R. Therefore, generating predictions about observable human behavior (e.g., reading times) from these parsers requires making some additional linking hypotheses. Most prior hypotheses that link parsing decisions to human behavior have focused on notions of processing effort, such as the number of parse states explored Hale (2011), the maximum number of items on the stack at any given point Joshi (1990), or the maximum amount of time a node stays in memory De Santo (2021). These hypotheses cannot be used to generate priming predictions because they do not specify a mechanism by which a prime sentence might facilitate the processing of a target sentence.
One notable exception is the ACT-R based left-corner repair parser proposed by Lewis and Vasishth (2005), in which parsing decisions are made based on the activation of different chunks in the memory (such as words or grammar rules). The activation of chunks in this model can capture notions of both processing difficulty and priming. However, this model assumes a strong dissociation between the grammar and the lexicon and therefore cannot be adopted directly to generate predictions from lexicalized grammar formalisms such as Minimalist Grammar Stabler (1996), Combinatorial grammar Steedman (1988), Lexical-Functional Grammar Kaplan et al. (1981) or Head-Driven Phrase Structure Grammar Pollard and Sag (1994). The goal of this work is to develop a framework that can be used to generate priming predictions from any theory of syntax that can generate parse trees.
@@ 2.3 Prior models of priming
Reitter et al. (2011) proposed an ACT-R based model of priming in which lexical and grammatical knowledge are more strongly interconnected than in the parser proposed by Lewis and Vasishth (2005). However, this model can only generate sentences given a semantic description, and therefore can only be used to model sentence production and not sentence processing. The models of priming that can model processing either do not explicitly model syntactic structure (Chang et al., 2006; Malhotra, 2009; Prasad et al., 2019; Sinclair et al., 2022) or do not explicitly implement the priming mechanism (Snider, 2008). Hence, the goal of this work is to bridge this gap and develop a model that explicitly models both syntactic structure and the priming mechanism, which can be used to generate behavioral processing predictions.
@@ 3 Model description
SPAWN uses the three components of ACT-R that are relevant for parsing: declarative memory (§ 3.1), which contains information about lexical and syntactic categories (cf. Reitter et al., 2011); procedural memory (§ 3.2), which contains the algorithm for retrieving syntactic categories from memory and combining them together; and buffers, which store the words the parser has encountered so far, the syntactic categories retrieved for those words and the current parse state.
@@ 3.1 Declarative memory (the grammar)
Declarative memory in SPAWN consists of two types of chunks (sets of attribute-value pairs): syntax chunks and lexical chunks.
Lexical chunks:
Each lexical chunk stores a lexical item in the vocabulary along with the set of syntactic categories that the word could be associated with. For example, the lexical chunk for the word “by” encodes that it needs to be associated with the syntactic category “preposition”, whereas the lexical chunk for “examined” encodes that it can be either be associated with the “transitive verb” category or the “past participle” category.
Syntax chunks:
Each syntax chunk stores the constraints on the contexts in which a category can occur. For example, the past participle category encodes that it needs to have a prepositional phrase category on its right; transitive verb encodes that it needs to have a determiner phrase category on its left and right.111The CCG notation to encode the “transitive verb” category is (TPDP)/DP; the forward slask indicates the words needs to combine with a DP on the right and the forward slash that it needs to combine with DP on the left. TP is the category that results from this combination. Since Combinatorial Categorial Grammar (CCG; Steedman, 1988) is designed to express such constraints, we use this grammar formalism to represent our syntax chunks.
@@ 3.2 Procedural memory (the parser)
SPAWN parses sentences incrementally, one token at a time. For each token, it retrieves a syntactic category (§ 3.2.1), and tries to integrate it with the current parse state (§ 3.2.2). If integration is not successful, then it retrieves another category and repeats the process. If none of the possible tags can successfully integrate with the current parse state, a reanalysis is triggered; depending on the model setting, the parser either goes back to the start of the sentence, or selects a previous token to reanalyze (§ 3.2.3). Once the integration is successful, the parser decides whether the next token is a null token (a token not associated with any overt word in the sentence) and if yes, adds it to the sequence of tokens (§ 3.2.4). Finally, the parser proceeds to the next token. This retrieval-integration-reanalysis-null prediction continues until all of the tokens in the sentence are assigned syntactic categories.
@@ 3.2.1 Retrieval
When processing a word , the parser retrieves the category with the highest activation from the set of all possible categories that can be associated with. The activation for any category is given by Equation 1, where is the base-level activation, is the activation spreads to , is the inhibition from the buffer to the when processing sentence , and is noise that is sampled from a specified normal distribution.
Base-level activation:
@@
The time to process a word is given by Equation 3, where is the number of chunks retrieved when processing , the activation of the -th chunk the model retrieved when processing (as computed using Equation 1), a latency factor and a latency exponent; these hyperparameters are set based on prior ACT-R models (see Appendix 5.3).
Lexical activation:
The context independent activation a word spreads to a category is given in Equation 4, where , another hyperparameter set based on prior ACT-R models (see Appendix 5.3), is the maximum activation that any word can spread.
Inhibition:
@@ 3.2.2 Integration
Integrating a retrieved syntactic category involves combining with the current parse state ; this combination is determined by the CCG composition process (Steedman, 1996). The parser starts by sequentially trying to apply each of the six composition rules (see Appendix B.2), stopping once a successful combination is found. If no successful combination is found, then the parser tries to apply each of the type-raising rules222A type-raising rule allows the incremental composition of categories. For example consider the category for transitive verb (TPDP)/DP. The parentheses dictate that the DP on the right of the verb needs to combine with the verb before the DP on the left. This is a problem because words are processed from left-to-right in English. A type-raising rule raises the DP on the left of the verb to TP/(TPDP), which lets it combine with the transitive verb category to result in the state TP/DP. in the grammar to the retrieved tag or parse state and then sequentially apply all six composition rules. If no successful combination is possible after type raising, then the retrieved category cannot be integrated into the current parse state; the parser then needs to either retrieve another category for the word being processed, or, if no unexplored categories remain, trigger a reanalysis.
@@ 3.2.3 Reanalysis
SPAWN uses an inhibition-based backtracking mechanism for reanalysis. When a reanalysis gets triggered at , the parser selects an index to backtrack to, where . We use two methods to select the reanalysis index. In one, we always select the first word of the sentence. In the other, we sample the index weighted by the number of categories that words at each index can be associated with (i.e., the amount of context-independent uncertainty at each index). Once it selects , it discards all of the categories retrieved for , and resets the parse state to what it was at . The parser keeps track of the categories that were discarded for all the words when processing a given sentence , and uses this to compute the inhibition for each category using Equation 5.333While storing all of the discarded categories is a convenient way to compute inhibition, it is not cognitively plausible. Future work can examine other ways of computing inhibition that relies on summaries of discarded categories, instead of storing all of the instances.
While the inhibition term decreases the probability of repeatedly retrieving the same incorrect category for a word every time it is reanalyzed, it does not eliminate it: inhibition might be insufficient to suppress the relative activation of an incorrect category , it has very high base-level or lexical activation. To ensure that the parsing process eventually results in the right parse, we implemented a “give-up” mechanism, where after iterations (set to ), the model uses only inhibition and noise to compute activation of .
3.2.4 Null or covert elements
SPAWN assumes that only certain parse states can be followed by null elements; this assumption ensures that the parser does not try to insert null elements after every token, which would result in an exponential increase in the search space because every null token could be followed by other null tokens. When the current parse state can be followed by a null element, there is uncertainty about the next parsing decision. Consider the following sentence, which is not related to relative clauses or part of our experimental setup, but illustrates the uncertainty in a theory-independent way.
. The cat examined the doctor and the doctor examined the cat.
@@ 3.2.4 Null or covert elements
This sentence is grammatical with or without the bolded red word. Therefore, after parsing the prefix “The cat examined the doctor and the doctor”, the parser has to predict whether the upcoming verb in the sentence is covert or overt. If it expects the next verb to be a covert “examined”, it should retrieve a null transitive verb category and integrate it with the current parse state before processing the remainder of the sentence. On the other hand, if it expects an overt “examined”, the parser should not retrieve any null elements — retrieving a null transitive verb would make the sentence ungrammatical when it encounters the overt transitive verb.
@@
We model this decision in SPAWN in the same way that we model other uncertainty: pick the option with the highest activation, where corresponds to the current word and , where correspond to the different types of null elements that can come after the current parse state. The activation for is given by Equation 6:
@@
is computed the same as lexical activation for syntactic categories (Equation 4), and the same as inhibition (Equation 5). As in Equation 1, is noise sampled from a normal distribution. We do not include base-level activation for the null categories in this computation, because the base-level activation for the not-null category would be extremely high — most sentences in the corpus do not have null elements —and would therefore dominate the overall activation, resulting in the null categories never being retrieved.
@@ 3.3 Updating activations (“learning”)
Learning in SPAWN occurs by updating the counts of syntactic categories, which in turn are used to compute base-level and lexical activations (Equations 2, 4). These counts are updated at the end of processing each sentence based on the final set of categories and null-elements that were retrieved.
3.3.1 How does priming emerge?
SPAWN does not explicitly model priming. Rather, priming emerges as a consequence of parsing and learning. There are two reasons priming emerges: an increase in the activation of relevant categories and an increase in the probability of reanalysis.
Increased activation:
[…]
. The lawyer sent the client the letter.
Let us assume that in 3.3.1, the verb “sent” is associated with a PO category, whereas in 3.3.1 it is associated with a DO category. Therefore the base-level activation for PO category, as well as the lexical activation from “sent” to PO will be higher in a model that has just parsed 3.3.1 than in a model that has just parsed 3.3.1. Now consider that a model which has processed one of these sentences is presented with a partial target sentence “The boy sent the ___”. The model is more likely to retrieve the PO category for “sent”, and thus more likely to produce a sentence like 3.3.1 if it has just processed 3.3.1, compared to if it has just processed 3.3.1.
. The boy sent the gift to his friend.
Increased probability of re-analysis:
Consider a model that just processed a sentence like 3.3.1, where the verb “sent” is assigned SHIFT category.
. The lawyer sent to his client the letter.
After processing this sentence, the model is more likely to expect a “to” after the verb “sent” than “the”. Therefore, when it encounters the partial target “The boy sent the” a reanalysis gets triggered. The probability of the model assigning PO category to “sent” is higher when it initially retrieves a SHIFT category and a reanalysis gets triggered, than when it initially retrieves a DO category. Therefore, SPAWN would predict that the probability of the target being completed with 3.3.1 is higher when it is preceded by 3.3.1 compared to 3.3.1.
4 A case study: competing theories of reduced relative clauses

### v2

@@ Abstract
Structural priming is a widely used psycholinguistic paradigm to study human sentence representations. In this work we introduce SPAWN, a cognitively motivated parser that can generate quantitative priming predictions from contemporary theories in syntax which assume a lexicalized grammar. By generating and testing priming predictions from competing theoretical accounts, we can infer which assumptions from syntactic theory are useful for characterizing the representations humans build when processing sentences. As a case study, we use SPAWN to generate priming predictions from two theories (Whiz-Deletion and Participial-Phase) which make different assumptions about the structure of English relative clauses. By modulating the reanalysis mechanism that the parser uses and strength of the parser’s prior knowledge, we generated nine sets of predictions from each of the two theories. Then, we tested these predictions using a novel web-based comprehension-to-production priming paradigm. We found that while the some of the predictions from the Participial-Phase theory aligned with human behavior, none of the predictions from the the Whiz-Deletion theory did, thus suggesting that the Participial-Phase theory might better characterize human relative clause representations.
@@ 1 Introduction
Prior work Branigan et al. (1995) found that targets like 1 were produced more often, and were processed more rapidly, when they were preceded by primes like 1, that have the same structure, than when they were preceded by primes like 1, which, while describing the same transfer event as 1, have a different structure.
@@
From this result, Branigan et al. inferred that participants’ mental representation of 1 is more similar to that of 1 than of 1.
Branigan and Pickering (2017) propose that by carefully studying which sentences prime each other we can build a theory of human structural representations. Building such a theory requires us to generate hypotheses about the particular prime-target pairs that would be most informative to compare. Insights from theoretical syntax, a field that has spent decades studying the structure of sentences, can help constrain this hypothesis space Gaston et al. (2017): if two theories generate different priming predictions, the theory whose prediction better aligns with human behavior better characterizes the representations humans build. In this work we introduce a new parser, the Serial Parser in ACT-R With Null elements (SPAWN), that can generate quantitative priming predictions from theories in syntax.
SPAWN is a cognitively motivated parser in which the parsing decisions are driven by the computational principles proposed by a general purpose cognitive architecture, Adaptive Control of Thought-Rational (ACT-R; Anderson et al., 2004). Thus, SPAWN not only describes the computations underlying human parsing, but also specifies the cognitive processes involved. This level of specification makes it possible to explain why, given a grammar, some sentence A is primed more by sentence B compared to C, which in turn is useful for generating quantitative behavioral priming predictions from syntactic theories.
Existing algorithmic models of parsing with this level of specification Lewis and Vasishth (2005) are limited in their ability to model assumptions from theories that use more contemporary frameworks like Minimalism for two reasons: First, they assume a disconnect between lexical and grammatical knowledge, which is inconsistent with the lexicalized grammar formalisms these frameworks adopt; Second, the models do not specify mechanisms to handle null (or covert) lexical items, which are essential components of several contemporary syntactic theories. SPAWN bridges this gap by adopting a lexicalized grammar formalism and specifying an explicit mechanism for null items.
@@
We generate priming predictions from two competing syntactic theories: Whiz-Deletion (Chomsky, 1965), which assumes that the structure of 1 is identical to the structure of 1, but that the words “which” and “was” are covert; and Participial-Phase (Harwood, 2018) which assumes that 1 and 1 have different structures. We describe these theories in more detail in § 4. We generate nine sets of predictions from the two theories by modulating two factors: First, the strength of prior knowledge (model exposed to 0, 100 or 1000 sentences before the experiment); Second, the reanalysis mechanism (model goes back to the beginning of the sentence, or model uses one of two entropy-based measures to select a word to go back to). Then, we compare the predictions from these two theories to empirical human data we collected using a novel web-based comprehension-to-production priming paradigm.
We found that the predictions from the Whiz-Deletion never aligned with the qualitative pattern of human priming behavior, whereas under some assumptions about the underlying reanalysis mechanism and strength of prior knowledge, predictions from the Participial-Phase theory did align with the qualitative empirical pattern. These results suggest that the Participial-Phase account better characterizes human sentence representations. More broadly, this case study highlights how SPAWN can be used to adjudicate between competing theoretical assumptions: the quantitative behavioral predictions SPAWN generates can clarify how differences in assumptions about sentence structure or parsing mechanisms might translate into testable behavioral differences (if at all).
@@ 2 Background
2.1 The ACT-R framework
ACT-R is a cognitive architecture designed to explain cognition through a small set of general computational principles and mechanisms that are relevant to a wide range of tasks and domains. One such mechanism which is particularly relevant in SPAWN is the retrieval of information from memory. The specific computational principles and algorithms that guide retrieval in ACT-R are outlined in § 3.2.1. Crucially, since ACT-R is intended to be a general purpose cognitive mechanism, most of the hyperparameters involved in this algorithm are already fixed based on data from a wide range of experimental paradigms and cognitive phenomena. This restricts the degrees of freedom and constrains the space of predictions that can be generated from any given theory.
@@ 2.2 Prior models of parsing
In most existing symbolic and neural-network based parsers, parsing decisions are not driven by specific cognitive principles such as the ones proposed by ACT-R. Therefore, generating predictions about observable human behavior (e.g., reading times) from these parsers requires making some additional linking hypotheses. Most prior hypotheses that link parsing decisions to human behavior have focused on notions of processing effort, such as the number of parse states explored Hale (2011), the maximum number of items on the stack at any given point Joshi (1990), or the maximum amount of time a node stays in memory De Santo (2021). These hypotheses cannot be used to generate priming predictions because they do not specify a mechanism by which a prime sentence might facilitate the processing of a target sentence.
One notable exception is the ACT-R based left-corner repair parser proposed by Lewis and Vasishth (2005), in which parsing decisions are made based on the activation of different chunks in the memory (such as words or grammar rules). The activation of chunks in this model can capture notions of both processing difficulty and priming. However, this model assumes a strong dissociation between the grammar and the lexicon and therefore cannot be adopted directly to generate predictions from lexicalized grammar formalisms such as Minimalist Grammar Stabler (1996), Combinatorial grammar Steedman (1988), Lexical-Functional Grammar Kaplan and Bresnan (1981) or Head-Driven Phrase Structure Grammar Pollard and Sag (1994). SPAWN is an ACT-R parser that models the link between the grammar and the lexicon can therefore generate predictions from lexicalized grammars.
@@ 2.3 Prior models of priming
While several models of priming have been proposed, as we illustrate in Figure 1, none of them can be used to adjudicate between contemporary syntactic theories. Many models of priming that model sentence processing either do not explicitly model syntactic structure (Chang et al., 2006; Malhotra, 2009; Prasad et al., 2019; Sinclair et al., 2022) or do not explicitly implement the mechanisms that result in priming (Snider, 2008). Reitter et al. (2011) proposed an ACT-R based model of priming that does explicitly implement priming mechanisms and, unlike Lewis and Vasishth’s ACT-R model, also assumes a strong link between lexical and grammatical knowledge, and is thus consistent with contemporary lexicalized grammar formalisms. However, this model can only generate sentences given a semantic description, and therefore can only be used to model sentence production and not sentence processing. We bridge this gap with SPAWN.
@@ 3 Model description
SPAWN uses the three components of ACT-R that are relevant for parsing: declarative memory, which contains information about lexical and syntactic categories (cf. Reitter et al., 2011); procedural memory, which contains the algorithm for retrieving syntactic categories from memory and combining them together; and buffers, which store the words the parser has encountered so far, the syntactic categories retrieved for those words and the current parse state.111https://github.com/grushaprasad/spawn We describe the two memory components below (§ 3.1, § 3.2), as well as the mechanisms for learning and priming (§ 3.3, § 3.4).
@@ 3.1 Declarative memory (the grammar)
Declarative memory in SPAWN consists of two types of chunks (sets of attribute-value pairs): syntax chunks and lexical chunks (see § A.3 for the entire list of syntax and lexical chunks we use in this work).
Lexical chunks
Each lexical chunk stores a word in the vocabulary along with the set of syntactic categories that the word could be associated with. For example, the lexical chunk for “examined” encodes that it can be either be associated with the transitive verb category or the past participle category.
Syntax chunks
Each syntax chunk stores the constraints on the contexts in which a category can occur. For example, the transitive verb category encodes that it needs to have a determiner phrase category on its left and right. We use the Combinatorial Categorial Grammar (CCG; Steedman, 1988) formalism to express such constraints.222The CCG notation to encode the “transitive verb” category is (TPDP)/DP; the forward slash indicates the words needs to combine with a DP on the right and the forward slash that it needs to combine with DP on the left. TP is the category that results from this combination.
@@ 3.2 Procedural memory (the parser)
SPAWN parses sentences incrementally, one word at a time. As schematized in Figure 2, processing each word involves four steps: retrieval, reanalysis, integration and null-prediction.
@@ 3.2.1 Retrieval
When processing a word in a sentence , the parser retrieves the category with the highest activation from the set of all possible categories that can be associated with. The activation for any category is given by Equation 1, where is the base-level activation, is the activation spreads to , is the inhibition from the buffer to the when processing sentence , and is noise sampled from .
Base-level activation
@@
The time to process a word is given by Equation 3, where is the number of chunks retrieved when processing , the activation of the -th chunk the model retrieved when processing (computed using Equation 1), a latency factor and a latency exponent.
Thus, in Equation 2 is .
Lexical activation
The context independent activation a word spreads to a category is given in Equation 4, where is the maximum activation that any word can spread.
Inhibition
@@ 3.2.2 Integration
Integrating a retrieved syntactic category involves combining with the current parse state ; this combination is determined by the CCG composition process (Steedman 1996; § B.1). If no successful combination is possible, then the retrieved category cannot be integrated into the current parse state; the parser then needs to either retrieve another category for the word, or, if no unexplored categories remain, trigger a reanalysis.
@@ 3.2.3 Reanalysis
When a reanalysis gets triggered at , the parser selects an index to regress to, where . The method used to select is a hyperparameter with two settings: first-word regression (go back to the first word every time) and entropy-weighted regression (sample from weighted by the parser’s uncertainty at each index). Once the parser selects , it discards all of the categories retrieved for , and resets the parse state to what it was at . The parser keeps track of the categories that were discarded when processing each word in a sentence , and uses this to compute the inhibition for each category using Equation 5.
Calculating uncertainty
To calculate uncertainty at index in entropy-weighted regression, we computed the activation of each category associated with by adding together and (§ 3.2.1). Then, we converted these activation values into probabilities with the softmax function (temperature 1 or 10), and finally computed the entropy from these probabilities.
“Give-up” mechanism
Despite inhibiting previously discarded categories, the parser could still get stuck in a loop retrieving the same (incorrect) category every time it is processing if has a very high base-level or lexical activation. To prevent an infinite loop, we implemented a “give-up” mechanism, where after iterations, the model ignores the base-level and lexical activation and uses only inhibition and noise to compute activation of . Setting to 100 or 1000 resulted in nearly identical results (§ D).
3.2.4 Null or covert element prediction
Null or covert elements in sentences add additional uncertainty to the parsing process. To illustrate this, let us consider an example that is unrelated to our experimental setup, but illustrates the uncertainty in a theory-independent way. Given a prefix “The cat examined the doctor and the doctor …” consider the following continuations; * indicates the continuation is ungrammatical.
. … examined the cat.
. … NULLexamined the cat.
. * … NULLexamined examined the cat.
@@ 3.2.4 Null or covert element prediction
The covert NULLexamined can only occur if its overt counterpart is absent. Therefore, after parsing the prefix, a serial parser has to predict whether the upcoming word in the sentence is covert or overt. If it expects the next word to be overt “examined”, the parser should not retrieve any null elements. On the other hand, if it expects the next word to be covert NULLexamined, it needs to retrieve this category and integrate it with the current parse state before processing the remainder of the sentence.
@@
We model this decision in SPAWN in the same way that we model other uncertainty: pick the option with the highest activation, where is the current word, and , where are the types of null elements that can come after the current parse state. The activation for is given by Equation 6:
@@
and are the same as in Equations 4 and 5. As in Equation 1, is noise sampled from . We do not include base-level activation for the null categories in this computation, because the base-level activation for the not-null category would be extremely high (most sentences in the corpus do not have null elements) and would result in the null categories never being retrieved. We also assume that only certain parse states can be followed by null elements (§ B.2): if the parser tried to insert null elements after every word, it would result in an exponential increase in the search space.
@@ 3.3 Updating activations (“learning”)
3.4 Emergence of priming in SPAWN
Priming in SPAWN emerges as a consequence of parsing and learning. There are two factors that can result in priming: an increase in the activation of relevant categories and an increase in the probability of reanalysis.
Increased activation
When a word in the target sentence is ambiguous between two categories and , if the parser retrieved in a preceding prime sentence, that increases its base and lexical activation relative to , which makes more likely to be retrieved in the target as well.
Increased reanalysis
When a word in the target sentence is ambiguous between two categories and , and has higher base and lexical activation, then the parser is more likely to retrieve initially. If a sequence of parsing decisions causes the parser to reanalyze the word, then the probability of the parser eventually retrieving increases: the inhibition to during reanalysis decreases the difference in activation between and .
4 A case study: Evaluating competing theories of reduced relative clauses